OpenStudy (anonymous):

These four lines have been graphed on the same coordinate grid. Which lines are parallel to each other?

OpenStudy (anonymous):

line j: y = 1/4x + 4 line k: y = 1/4x + 3 line l: y = 3x + 4 line m: y = 4x + 1/4

OpenStudy (phi):

the ones that have the same slope. Slope is the number in front of the x

OpenStudy (anonymous):

so, j and k?

OpenStudy (phi):

yes 1/4 equals 1/4 so that is the same number. those lines are parallel

OpenStudy (anonymous):

thanks so much! i appreciate it! :D